OpenStudy (anonymous):

A pair of equations is shown below: y = 2x − 1 y = 4x − 5 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations. Part B: What is the solution to the pair of equations? Help explain how to do this

OpenStudy (anonymous):

If the RHS of both equations equal y, then they have to equal each other. Solve the following for x. 4 x-5=2 x-1 2x=4 x=2 y =2(2) - 1 y = 3